# Category $\mathcal{O}$ for hybrid quantum groups and non-commutative Springer resolutions

### v3

@@ 1.1. Hybrid quantum group and its category ${\mathcal{O}}$
Let $G$ be a complex connected and simply-connected semisimple algebraic group, with a Borel subgroup $B$ and a Cartan subgroup $T\subset B$. Let $U_{q}$ be Lusztig’s quantum group and let ${\mathfrak{U}}_{q}$ be De Concini–Kac’s quantum group, associated to $G$. The hybrid (or mixed) quantum group $U^{\mathrm{hb}}_{q}$ is an algebra admitting a triangular decomposition  [^1] In the present paper we will use the definition $U^{\mathrm{hb}}_{q}={\mathfrak{U}}^{-}_{q}\otimes{\mathfrak{U}}^{0}_{q}\otimes U^{+}_{q}$, for the convenience of deforming category ${\mathcal{O}}$. In fact these two definitions lead to the same category ${\mathcal{O}}$ (see [9, Lem 3.1]).
$U^{\mathrm{hb}}_{q}={\mathfrak{U}}^{-}_{q}\otimes U^{0}_{q}\otimes U^{+}_{q},$
which was firstly introduced by Gaitsgory [22] with the perspective of generalizing the Kazhdan–Lusztig equivalence. It also appears naturally when one considers the quantum analogue of the BGG category ${\mathcal{O}}$. Recall that the BGG category ${\mathcal{O}}$ (with integral weights) for the Lie algebra ${\mathfrak{g}}=\mathrm{Lie}(G)$ can be interpreted as the category
@@
of finitely generated ${\mathfrak{g}}$-modules whose restriction on ${\mathfrak{b}}=\mathrm{Lie}(B)$ is integrable. Therefore, one can define the quantum category ${\mathcal{O}}$ to be the category
@@
of finitely generated ${\mathfrak{U}}_{q}$-modules equipped with an integrable $U^{\geq}_{q}$-module structure that is compatible through ${\mathfrak{U}}^{\geq}_{q}\rightarrow U^{\geq}_{q}$. It is exactly the category ${\mathcal{O}}$ associated to $U^{\mathrm{hb}}_{q}$ with its triangular decomposition.
@@ 1.2. Main result
Let ${\mathrm{A}}$ be the ($G$-equivariant) ${\mathbb{C}}[{\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]$-algebra defined in loc. cit. as the non-commutative counterpart of the Grothendieck resolution $\tilde{{\mathfrak{g}}}=G\times^{B}{\mathfrak{b}}\rightarrow{\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}$ (where ${\mathfrak{t}}=\mathrm{Lie}(T)$ and $W$ is the Weyl group). Let $\pi_{{}_{\tilde{{\mathcal{N}}}}}:\tilde{{\mathcal{N}}}=G\times^{B}{\mathfrak{n}}\rightarrow{\mathfrak{g}}$ be the Springer resolution (where ${\mathfrak{n}}$ is the nilpotent radical of ${\mathfrak{b}}$). Our main result is
Theorem A (=Corollaries 5.15, 6.8).
@@
(1.1) ${\mathcal{O}}^{0}_{\zeta}{\ \xrightarrow{\sim}\ }\mathrm{Coh}^{G}(\pi_{{}_{\tilde{{\mathcal{N}}}}}^{*}{\mathrm{A}})$
@@
Note that a similar result for modular category ${\mathcal{O}}$ was obtained by Losev [26] recently. The category $\mathrm{Coh}^{G}(\pi_{{}_{\tilde{{\mathcal{N}}}}}^{*}{\mathrm{A}})$ firstly appeared in [10], which was shown to be derived equivalent to a version of the affine Hecke category and coincide with the heart of the “new t-structure” defined by Frenkel–Gaitsgory [21]. It would be interesting to explore the relationship between (1.1) and a conjectural equivalence of Gaitsgory [22, Conj. 9.2.2] recently proved by Chen–Fu [16].
@@ 1.3. Equivalence of Steinberg block and functor ${\mathbb{V}}$
(1.2) ${\mathcal{O}}^{-\rho}_{\zeta}{\ \xrightarrow{\sim}\ }\mathrm{Coh}^{G}(\tilde{{\mathcal{N}}}).$
Based on the equivalence above, we establish functors ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$ on both sides of (1.1) that are fully-faithful on the objects admitting Verma flags (in the deformed setting, see §1.5 below).
@@
${\mathbb{V}}_{c}:\ \mathrm{Coh}^{G}(\pi_{{}_{\tilde{{\mathcal{N}}}}}^{*}{\mathrm{A}})\rightarrow\mathrm{Coh}^{G}(\tilde{{\mathcal{N}}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})$
by base change along $\pi_{{}_{\tilde{{\mathcal{N}}}}}:\tilde{{\mathcal{N}}}\rightarrow{\mathfrak{g}}$ of the non-commutative resolution ${\mathrm{A}}\text{-}\mathrm{mod}\rightarrow\mathrm{Coh}({\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})$. To relate the images of two functors ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$, we use a $C$-linear extension of the equivalence (1.2), i.e.
${\mathcal{O}}^{-\rho}_{\zeta,C}{\ \xrightarrow{\sim}\ }\mathrm{Coh}^{G}(\tilde{{\mathcal{N}}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}).$
Then we compare the images of the projective objects under the functors ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$. To that end, we firstly establish a “de-equvariantization” of Theorem A.
@@ 1.4. A de-equvariantization
More precisely, let ${\mathfrak{U}}^{b}_{\zeta}$ be the image of ${\mathfrak{U}}_{\zeta}\rightarrow U^{\mathrm{hb}}_{\zeta}$, and consider the category ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{T}$ of $X^{*}(T)$-graded representations of ${\mathfrak{U}}^{b}_{\zeta}$. Let ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{T,0}$ be its principal block. Note that the restriction of $\pi_{{}_{\tilde{{\mathcal{N}}}}}^{*}{\mathrm{A}}$ on the fiber ${\mathfrak{n}}\hookrightarrow\tilde{{\mathcal{N}}}$ is ${\mathrm{A}}\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}[{\mathfrak{n}}]$. Hence we can identify
$\mathrm{Coh}^{G}(\pi_{{}_{\tilde{{\mathcal{N}}}}}^{*}{\mathrm{A}})=\mathrm{Coh}^{B}({\mathrm{A}}\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}[{\mathfrak{n}}])$
@@
(1.3) ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{T,0}{\ \xrightarrow{\sim}\ }\mathrm{Coh}^{T}({\mathrm{A}}\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}[{\mathfrak{n}}]),$
@@
The proof of Proposition 1.2 is based on results of Tanisaki [42], where the localization theorem for ${\mathfrak{U}}_{\zeta}$ was established and it further related ${\mathfrak{U}}_{\zeta}$-modules to ${\mathrm{A}}$-modules by some splitting bundles for the sheaf of differential operators on the quantized flag variety. Note that it is not clear from their constructions whether the splitting bundles are $U_{\zeta}$-equivariant or not. So à priori, one can not derive (1.1) from (1.3), by adding the $U^{\geq}_{\zeta}$-equivariant and $B$-equivariant conditions on both sides. However, after proving our main result, we can show that (1.3) dose respect the $B$-actions on both categories as a consequence (see Corollary 6.7).
@@ 1.5. Deformations
For the classical category ${\mathcal{O}}$, the deformation technique played an important role on studying basic structures, e.g. [38], [40] and [20]. Deformation of ${\mathcal{O}}_{\zeta}$ was studied in [37]. Let $S={\mathbb{C}}[{\mathfrak{t}}]_{\hat{0}}$ be the completion of ${\mathbb{C}}[{\mathfrak{t}}]$ at $0\in{\mathfrak{t}}$. The deformation category for ${\mathcal{O}}_{\zeta}$ is an $S$-linear category ${\mathcal{O}}_{\zeta,S}$ whose specialization to the residue field is ${\mathcal{O}}_{\zeta}$. Let ${\mathrm{A}}_{({\mathfrak{t}}/W)_{\hat{0}}}$ be the completion of ${\mathrm{A}}$ at $0\in{\mathfrak{t}}/W$. We have a deformed version of Theorem A
@@
(1.4) ${\mathcal{O}}^{0}_{\zeta,S}{\ \xrightarrow{\sim}\ }\mathrm{Coh}^{G}(\pi^{*}{\mathrm{A}}_{({\mathfrak{t}}/W)_{\hat{0}}}),$
where ${\mathcal{O}}^{0}_{\zeta,S}$ is the principal block of ${\mathcal{O}}_{\zeta,S}$ and $\pi:\tilde{{\mathfrak{g}}}\rightarrow{\mathfrak{g}}$ is the natural projection.
@@ 1.6. Kazhdan–Lusztig polynomial
Consider the ${\mathbb{C}}^{\times}$-action on ${\mathfrak{g}}$ by $t.x=t^{-2}x$, for any $t\in{\mathbb{C}}^{\times}$, $x\in{\mathfrak{g}}$. The ${\mathbb{C}}[{\mathfrak{g}}]$-algebra ${\mathrm{A}}$ can be equipped with a ${\mathbb{C}}^{\times}$-equivariant structure. By the equivalence (1.1), we can define a grading on ${\mathcal{O}}^{0}_{\zeta}$ via
${\mathcal{O}}^{0,{\mathrm{gr}}}_{\zeta}:=\mathrm{Coh}^{G\times{\mathbb{C}}^{\times}}(\pi_{{}_{\tilde{{\mathcal{N}}}}}^{*}{\mathrm{A}})\ \rightarrow\ {\mathcal{O}}^{0}_{\zeta}.$
@@ 1.7. Organization of the paper
In Section 4 we prove Proposition 1.2, and construct the functors ${\mathbb{V}}_{r}^{b}$, ${\mathbb{V}}^{b}_{c}$ in §4.2, whose fully-faithfulness will be proved in §4.4. In §4.2 we also study the induction functors from the categories in (1.3) to the categories in (1.1). In §4.3 we investigate the image of Verma modules under the equivalence (1.3).
In Section 5, we define the functors ${\mathbb{V}}_{r}$, ${\mathbb{V}}_{c}$, and show their compatibility with reflection functors. §5.2 is a slightly technical section. The projective objects in ${\mathcal{O}}_{\zeta}$ only exist after truncations, hence we have to define corresponding truncations on the coherent side and study projective objects in them, which will be done in §5.2. In §5.3 we show the fully-faithfulness of ${\mathbb{V}}_{r}$, ${\mathbb{V}}_{c}$ and then prove the main theorem.
In Section 6, we firstly recall the formalism of a rational group action on a category, then equip a $B$-action on LHS of (1.3) using coinduction functor. We apply our main theorem to show that the equivalence (1.3) respects $B$-actions on both categories, and the equivalence (1.1) can be obtained as the $B$-equivariantization of (1.3).
In Section 7, we firstly recall the periodic and generic polynomials. Then we discuss the liftings of the simple, Verma and projective objects in the graded principal block and compute the related graded multiplicities.
@@ 1.8. Conventions
$\mathrm{FN}_{X}(Y):=X_{\hat{z}}:=X\times_{\mathrm{Spec}R}\mathrm{Spec}(R_{\hat{z}}).$
@@ 2.1. Root data
Let $(X^{*}(T),X_{*}(T),\Phi,\check{\Phi})$ be the root datum associated with $G$. Let $\langle$-,$\rangle$ - , - ⟩ be the natural pairing of $\Lambda:=X^{*}(T)$ and $X_{*}(T)$. Let $\Phi^{+}\subset\Phi$ be the subsets of positive roots. Denote by $\alpha_{s}\in\Phi^{+}$ the simple root associated to $s\in{\mathrm{I}}$. Let ${\mathrm{Q}}\subset\Lambda$ be the root lattices. There is a partial order $\leq$ on $\Lambda$ given by $\mu\leq\lambda$ if $\lambda-\mu\in\sum_{s\in{\mathrm{I}}}{\mathbb{N}}\alpha_{s}$. Let $(d_{s})_{s\in{\mathrm{I}}}\in{\mathbb{N}}^{\mathrm{I}}$ be the unique tuple of relatively prime positive integers such that the matrix $\big{(}d_{s}\langle\check{\alpha}_{s},\alpha_{t}\rangle\big{)}_{s,t\in{\mathrm{I}}}$ is symmetry and positive definite. It defines a pairing $(-,-):{\mathrm{Q}}\times{\mathrm{Q}}\rightarrow{\mathbb{Z}}$ by $(\alpha_{s},\alpha_{t}):=d_{s}\langle\check{\alpha}_{s},\alpha_{t}\rangle$ and extents to

### v4

@@ 1.1. Hybrid quantum group and its category $\mathcal{O}$}
Let $G$ be a complex connected and simply-connected semisimple algebraic group, with a Borel subgroup $B$ and a Cartan subgroup $T\subset B$. Let $U_{q}$ be Lusztig’s quantum group [29] and let ${\mathfrak{U}}_{q}$ be De Concini–Kac’s quantum group [18], associated to $G$. The hybrid (or mixed) quantum group $U^{\mathrm{hb}}_{q}$ is an algebra admitting a triangular decomposition
$U^{\mathrm{hb}}_{q}={\mathfrak{U}}^{-}_{q}\otimes{\mathfrak{U}}^{0}_{q}\otimes U^{+}_{q}$,
which was firstly introduced by Gaitsgory [22] with the perspective of generalizing the Kazhdan–Lusztig equivalence. It also appears naturally when one considers the quantum analogue of the BGG category $\mathcal{O}$} [6]. Recall that the BGG category $\mathcal{O}$} (with integral weights) for the Lie algebra ${\mathfrak{g}}=\mathrm{Lie}(G)$ can be interpreted as the category
@@
of $B$-equivariant finitely generated $\mathfrak{g}$}-modules on which the actions of ${\mathfrak{b}}=\mathrm{Lie}(B)$ by restriction of $\mathfrak{g}$}-action and by differential of $B$-action coincide. Therefore, one can define the quantum category $\mathcal{O}$} to be the category
@@
of finitely generated ${\mathfrak{U}}_{q}$-modules equipped with an equivariant and integrable $U^{\geq}_{q}$-module structure that is compatible through ${\mathfrak{U}}^{\geq}_{q}\rightarrow U^{\geq}_{q}$. It is exactly the category $\mathcal{O}$} (with integral weights) associated to $U^{\mathrm{hb}}_{q}$ with its triangular decomposition.
@@ 1.2. Main result
Let $\mathsf{A}$} be the ($G$-equivariant) ${\mathbb{C}}[{\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]$-algebra defined in loc. cit. as the non-commutative counterpart of the Grothendieck resolution $\widetilde{{\mathfrak{g}}}=G\times^{B}{\mathfrak{b}}\rightarrow{\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}$ (where ${\mathfrak{t}}=\mathrm{Lie}(T)$ and $W$ is the Weyl group). Let $\pi_{{}_{\widetilde{{\mathcal{N}}}}}:\widetilde{{\mathcal{N}}}=G\times^{B}{\mathfrak{n}}\rightarrow{\mathfrak{g}}$ be the Springer resolution (where $\mathfrak{n}$} is the nilpotent radical of $\mathfrak{b}$}). Our main result is
Theorem A (=Corollaries 5.25&6.8).
@@
(1.1) ${\mathcal{O}}^{0}_{\zeta}{\ \xrightarrow{\sim}\ }\mathrm{Coh}^{G}(\pi_{{}_{\widetilde{{\mathcal{N}}}}}^{*}{\mathsf{A}})$
@@
Note that a similar result for modular analogue of BGG category $\mathcal{O}$} was obtained by Losev [25] recently. The category $\mathrm{Coh}^{G}(\pi_{{}_{\widetilde{{\mathcal{N}}}}}^{*}{\mathsf{A}})$ firstly appeared in [8], which was shown to be derived equivalent to a version of the affine Hecke category and coincide with the heart of the “new t-structure” defined by Frenkel–Gaitsgory [21]. It would be interesting to explore the relationship between (1.1) and a conjectural equivalence of Gaitsgory [22, Conj. 9.2.2] recently proved by Chen–Fu [16].
@@ 1.3. Equivalence of Steinberg block and functor $\mathbb{V}$}
[…]
Based on the equivalence above, we establish functors ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$ (here “rr” and “cc” represent “representation” and “coherent”, respectively) on both sides of (1.1) that are fully-faithful on the objects admitting Verma flags (in fact our constructions are in the deformed setting, see §1.5 below).
@@
${\mathbb{V}}_{c}:\ \mathrm{Coh}^{G}(\pi_{{}_{\widetilde{{\mathcal{N}}}}}^{*}{\mathsf{A}})\rightarrow\mathrm{Coh}^{G}(\widetilde{{\mathcal{N}}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})$
by base change along $\pi_{{}_{\widetilde{{\mathcal{N}}}}}:\widetilde{{\mathcal{N}}}\rightarrow{\mathfrak{g}}$ of the non-commutative resolution ${\mathsf{A}}\text{-}\mathrm{mod}\rightarrow\mathrm{Coh}({\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})$. To relate the images of two functors ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$, we use a $C$-linear extension of the equivalence (1.2), i.e.
${\mathcal{O}}^{-\rho}_{\zeta,C}{\ \xrightarrow{\sim}\ }\mathrm{Coh}^{G}(\widetilde{{\mathcal{N}}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}).$
Our next step is to compare the images of the projective objects under the functors ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$. To that end, we firstly establish a “de-equvariantization” of Theorem A.
@@ 1.4. A de-equvariantization
More precisely, let ${\mathfrak{U}}^{b}_{\zeta}$ be the image of ${\mathfrak{U}}_{\zeta}\rightarrow U^{\mathrm{hb}}_{\zeta}$, and consider the category ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{T}$ of $X^{*}(T)$-graded representations of ${\mathfrak{U}}^{b}_{\zeta}$. Let ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{T,0}$ be its principal block. Note that the restriction of $\pi_{{}_{\widetilde{{\mathcal{N}}}}}^{*}{\mathsf{A}}$ on the fiber ${\mathfrak{n}}\hookrightarrow\widetilde{{\mathcal{N}}}$ is ${\mathsf{A}}\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}[{\mathfrak{n}}]$. Hence we can identify
$\mathrm{Coh}^{G}(\pi_{{}_{\widetilde{{\mathcal{N}}}}}^{*}{\mathsf{A}})=\mathrm{Coh}^{B}({\mathsf{A}}\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}[{\mathfrak{n}}])$
@@
(1.3) ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{T,0}{\ \xrightarrow{\sim}\ }\mathrm{Coh}^{T}({\mathsf{A}}\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}[{\mathfrak{n}}]),$
@@
The proof of Proposition 1.2 is based on results of Tanisaki [39], where the localization theorem for ${\mathfrak{U}}_{\zeta}$ was established and it further related ${\mathfrak{U}}_{\zeta}$-modules to $\mathsf{A}$}-modules by some splitting bundles for the sheaf of differential operators on the quantized flag variety. Note that it is not clear from their constructions whether the splitting bundles are $U_{\zeta}$-equivariant or not. So à priori, one can not derive (1.1) from (1.3), by adding the $U^{\geq}_{\zeta}$-equivariant and $B$-equivariant conditions on both sides. However, after proving our main result, we can show that (1.3) does respect the $B$-actions on both categories as a consequence (see Corollary 6.7).
@@ 1.5. Deformations
For the classical category $\mathcal{O}$}, the deformation technique played an important role on studying basic structures, e.g. [36], [38] and [20]. Deformation of ${\mathcal{O}}_{\zeta}$ was studied in [35]. Let $S={\mathbb{C}}[{\mathfrak{t}}]_{\hat{0}}$ be the completion of ${\mathbb{C}}[{\mathfrak{t}}]$ at $0\in{\mathfrak{t}}$. The deformation category for ${\mathcal{O}}_{\zeta}$ is an $S$-linear category ${\mathcal{O}}_{\zeta,S}$ whose specialization to the residue field is ${\mathcal{O}}_{\zeta}$. Let $^{\mathsf{A}}_{({\mathfrak{t}}/W)_{\hat{0}}}$ be the completion of $\mathsf{A}$} at $0\in{\mathfrak{t}}/W$. We have a deformed version of Theorem A
@@
[…]
where ${\mathcal{O}}^{0}_{\zeta,S}$ is the principal block of ${\mathcal{O}}_{\zeta,S}$ and $\pi:\widetilde{{\mathfrak{g}}}\rightarrow{\mathfrak{g}}$ is the natural projection.
@@ 1.6. Kazhdan–Lusztig polynomial
[…]
${\mathcal{O}}^{0,{\mathrm{gr}}}_{\zeta}:=\mathrm{Coh}^{G\times{\mathbb{C}}^{\times}}(\pi_{{}_{\widetilde{{\mathcal{N}}}}}^{*}{\mathsf{A}})\ \rightarrow\ {\mathcal{O}}^{0}_{\zeta}.$
@@ 1.7. Organization of the paper
In Section 4 we prove the de-equivariantization of main result (i.e. Proposition 1.2) in §4.1 and its analogue for Steinberg block in §4.2. In §4.3 we study the induction functors from the categories in (1.3) to the categories in (1.1). We construct the functors ${\mathbb{V}}_{r}^{b}$, ${\mathbb{V}}^{b}_{c}$ in §4.4, whose fully-faithfulness will be proved in §4.6. In §4.5 we investigate the image of Verma modules under the equivalence (1.3).
In Section 5, we define the functors ${\mathbb{V}}_{r}$, ${\mathbb{V}}_{c}$, and show their compatibility with reflection functors in §5.1. The subsection §5.2 is slightly technical: we consider certain truncated subcategories of ${\mathcal{O}}^{0}_{\zeta}$ and of the coherent side, which admit enough projective objects and are compatible with functors ${\mathbb{V}}_{r}$, ${\mathbb{V}}_{c}$. In §5.3 we show the fully-faithfulness of ${\mathbb{V}}_{r}$, ${\mathbb{V}}_{c}$ and then prove the main theorem.
In Section 6, we firstly recall the formalism of a rational group action on a category in §6.1, then equip a $B$-action on LHS of (1.3) using coinduction functor in §6.2. In §6.3, we apply our main theorem to show that equivalence (1.1) is the $B$-equivariantization of (1.3), and obtain the compatibility of (1.1) with reflection functors as an application.
In Section 7, we firstly recall the periodic and generic polynomials in §7.1. Then we discuss the liftings of the simple, Verma and projective objects in the graded principal block and compute the related graded multiplicities in §7.2.
@@ 1.8. Conventions
[…]
@@ 2.1. Root data
Let $(X^{*}(T),X_{*}(T),\Phi,\check{\Phi})$ be the root datum associated with $G$. Let $\langle-,$-$\rangle$ be the natural pairing of $\Lambda:=X^{*}(T)$ and $X_{*}(T)$. Let $\Phi^{+}\subset\Phi$ be the subsets of positive roots. Denote by $\alpha_{s}\in\Phi^{+}$ the simple root associated to $s\in{\mathrm{I}}$. Let ${\mathrm{Q}}\subset\Lambda$ be the root lattices. There is a partial order $\leq$ on $\Lambda$ given by $\mu\leq\lambda$ if $\lambda-\mu\in\sum_{s\in{\mathrm{I}}}{\mathbb{N}}\alpha_{s}$. Let $(d_{s})_{s\in{\mathrm{I}}}\in{\mathbb{N}}^{\mathrm{I}}$ be the unique tuple of relatively prime positive integers such that the matrix $\big(d_{s}\langle\check{\alpha}_{s},\alpha_{t}\rangle\big)_{s,t\in{\mathrm{I}}}$ is symmetry and positive definite. It defines a pairing $(-,-):{\mathrm{Q}}\times{\mathrm{Q}}\rightarrow{\mathbb{Z}}$ by $(\alpha_{s},\alpha_{t}):=d_{s}\langle\check{\alpha}_{s},\alpha_{t}\rangle$ and extents to